Define a function A: N x N -> N as follows: A(m, n) ={2n, if m= 0; 0, if m≥1 and n= 0; 2, if m≥1 and n= 1; A(m-1, A(m, n-1)), if m≥1 and n≥2
(a) Calculate the following:
(i)A(1,0)
(ii)A(0,1)
(iii)A(1,1)
(iv)A(2,2).
1
Expert's answer
2020-12-15T19:44:29-0500

Consider the function "A:\mathbb N\times\mathbb N\to\mathbb N,"


"A(m,n)=\begin{cases}
2n,\ \ \ m=0\\
0,\ \ \ \ \ m\ge 1\text{ and } n=0\\
2, \ \ \ \ \ m\ge 1\text{ and } n =1\\
A(m-1,A(m,n-1)),\ \ \ \ m\ge 1\text{ and } n\ge 2
\end{cases}"


(i) Since "m=1" and "n=0" we use the formula "A(m,n)=0", and have that "A(1,0)=0"


(ii) Since "m=0" we use the formula "A(m,n)=2n", and have that "A(0,1)=2"


(iii) Since "m=1" and "n=1" we use the formula "A(m,n)=2", and have that "A(1,1)=2"


(iv) We use the formula "A(m,n)=2" for "m=2" and "n=1" to conclude that "A(2,1)=2".

In the following we use the formula "A(m,n)=A(m-1,A(m,n-1))" to calculate "A(2,2)" and "A(1,2)", and the formula "A(m,n)=2n" to calculate "A(0,2):"


"A(2,2)=A(1, A(2,1))=A(1,2)=A(0, A(1,1))=A(0,2)=4"